OpenStudy (boothepest1):

A rectangle has an area of 25cm2. If each side has its length doubled. What is the new area?

OpenStudy (mrnood):

area of original = l*h after doubling the sides they become 2l and 2h so new area is 2*l * 2*h = 4 *l*h = 4 times the original area